OpenStudy (anonymous):

substract: (8w-1/w^2-6w-16)-(w-6/w^2-4)

OpenStudy (anonymous):

please write your question clearly

OpenStudy (anonymous):

they are fractions

OpenStudy (anonymous):

abarnzy, This is from your problem statement: \[\left(8 w-\frac{1}{w^2}-6 w-16\right)-\left(w-\frac{6}{w^2}-4\right) \] OK?